OpenStudy (yacoub1993):

Simplify using properties of exponents. 16x234x14 A. 4x5/12 B. 4x1/12 C. 4x5/3 D. 12x1/12

OpenStudy (yacoub1993):

\[\frac{ 16x \frac{ 2 }{ 3 } }{4x \frac{ 1 }{ 4 } }\]

OpenStudy (yacoub1993):

I I need some help. I am really getting frustrated

OpenStudy (anonymous):

ok the translates into\[\frac{ 16 }{ 4 }*x ^{\frac{ 2 }{3 }-\frac{ 1 }{4 }}\]

OpenStudy (yacoub1993):

i think the equation is like this but not sure\[\frac{ 16 \times ^{\frac{ 2 }{3 }} }{ 4 \times ^{\frac{ 1 }{ 4 }}}\]

OpenStudy (yacoub1993):

let me snap a photo of the question @donny471

OpenStudy (anonymous):

\[4*x ^{\frac{ 8 }{ 12 }-\frac{ 3 }{ 12 }}=4x ^{\frac{ 5 }{ 12 }}\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Is this flvs?

OpenStudy (yacoub1993):

no its just one pic but its not uploading wen i try to upload it @Johnjakile1998

OpenStudy (anonymous):

Okay.. Can you click on the equation button and do it that way? But I think the answer is A, just from looking at the question. What type of math is this?

OpenStudy (yacoub1993):

Pre Calculus i am not sure whether it is (times) or (x)

OpenStudy (anonymous):

Its definitely (x). Not (times)

OpenStudy (anonymous):

I think the answer is A. I'm in Calculus, so you can trust me.

OpenStudy (anonymous):

@Yacoub1993

OpenStudy (yacoub1993):

thnxs @Johnjakile1998